\begin{thebibliography}{10}

\bibitem{AS} J.M. Anderson, A.L. Shields, Coefficient multipliers of Bloch functions, Trans. Amer. Math. Soc. 224 (1976), 255-265.

\bibitem{BST} G. Bennet, D.A. Stegenga, R.M. Timoney, Coefficients of Bloch and Lipschitz functions, Illinois J. Math. 25 (1981), 520-531.

\bibitem{BBG} K.D.\ Bierstedt, J.\ Bonet, A.\ Galbis, Weighted spaces
	of holomorphic functions on bounded domains, Michigan
Math.\ J.\ 40 (1993), 271--297.

\bibitem{BBT} K.D.\ Bierstedt, J.\ Bonet, J.\ Taskinen, Associated
	weights and spaces of holomorphic functions, Studia Math.\
	127 (1998), 137--168.

\bibitem{Bl} O. Blasco, Multipliers on spaces of analytic functions,
Canad. J. Math. 47 (1995), 44-64.

\bibitem{BG} O. Blasco, A. Galbis, On Taylor coefficients of entire functions integrable against exponential weights, Math. Nachr. 223 (2001), 5-21.

\bibitem{BP} O. Blasco, M. Pavlovic, Coefficient multipliers on Banach spaces of analytic functions, Rev. Mat. Iberoam. 27 (2011),  415-447.

\bibitem{BlZ} O. Blasco, C. Zaragoza-Benzosa, Multipliers on generalized
mixed norm sequence spaces, Abstr. Appl. Anal.  2014, Art. ID 983273

\bibitem{BDLT} J. Bonet, P. Doma\'nski, M. Lindstr\"om, J. Taskinen, Composition operators between weighted Banach spaces of analytic functions. J. Austral. Math. Soc. Ser. A 64 (1998), 101-118.

\bibitem{BT} J. Bonet, J. Taskinen, A note about Volterra operators on weighted Banach spaces of entire functions, Math. Nachr. 288 (2015), 1216-1225.

\bibitem{BRV} S.M. Buckley, M.S. Ramanujan, D. Vukotic, Bounded and compact multipliersbetween Bergman and Hardy spaces, Integral Equ. Oper. Theory 35 (1999), 1-19.

\bibitem{D} E. Doubtsov, Characterization of Hardy growth spaces with doubling weights, Bull. Austral. Math. Soc. 90 (2014), 275-282.

\bibitem{JP} M. Jevtic, M. Pavlovic, On the solid hull of the Hardy space, Michigan Math. J. 111 (2006), 439-446.

\bibitem{K} C. N. Kellogg, An extension of the Hausdorff-Young theorem,
Michigan Math. J. 18 (1971), 121-127.

\bibitem{Ki}  S.V. Kislyakov, S. V. Fourier coefficients of boundary values of functions that are analytic in the disc and bidisc. Spectral theory of functions and operators, II. Trudy Mat. Inst. Steklov. 155 (1981), 77–94, 183–185.

\bibitem{Li} H. Li, Convergence of Taylor series in Fock spaces, Studia Math. 220 (2014), 179-186.

\bibitem{L0} W. Lusky, On weighted spaces of harmonic and holomorphic functions, J. London Math. Soc. 51 (1995), 309-320.

\bibitem{L} W. Lusky, On the isomorphism classes of weighted spaces
of harmonic and holomorphic functions, Studia Math. 175 (1) (2006), 19-40.

\bibitem{SW} A.L. Shields, D.L. Williams, Bounded projections, duality
	and multipliers in spaces of analytic functions, Trans.
Amer. Math. Soc. 162 (1971), 287-302.

\bibitem{Tung} J. Tung, Taylor coefficients of functions in Fock spaces, J. Math. Anal. Appl. 318 (2006), 397-409.

\bibitem{V} D. Vukotic, On the coefficient multipliers of Bergaman spaces, J. London Math. Soc. 50 (1994), 341-348.


\end{thebibliography}

\documentclass[12pt]{article}
\usepackage{amsmath}
\usepackage{latexsym}
\usepackage{amsfonts}
\usepackage{amssymb}
\usepackage{amscd}
\usepackage{theorem}

\usepackage{color}

\newtheorem{theorem}{Theorem}[section]

\newtheorem{lemma}[theorem]{Lemma}
\newtheorem{proposition}[theorem]{Proposition}
\newtheorem{corollary}[theorem]{Corollary}

{\theorembodyfont{\rmfamily}

  \newtheorem{remark}[theorem]{Remark}
}

\newenvironment{proof}{    
  \noindent
  \textbf{Proof.}}{
  \hfill $\Box$
  \vspace{3mm}
}

\numberwithin{equation}{section}

\newcommand{\N}{\mathbb{N}} 
\newcommand{\R}{\mathbb{R}} 
\newcommand{\C}{\mathbb{C}} 
\newcommand{\D}{\mathbb{D}} 

\begin{document}

\title{Solid hulls of weighted Banach spaces of entire functions.  }

\author{Jos\'{e} Bonet  and Jari Taskinen }

\date{}

\maketitle

\begin{abstract}
Given a continuous, radial, rapidly decreasing  weight $v$ on the complex plane, we study the solid hull of its associated weighted space $H_v^\infty(\C)$ of all the entire functions $f$ such that $v|f|$ is bounded. The solid hull is found for a large
class of weights satisfying the condition (B) of Lusky. Precise formulations are
obtained for weights of the form $v(r)=\exp(-ar^p), a>0, p>0$. Applications to spaces of multipliers are included. 
\end{abstract}

\renewcommand{\thefootnote}{}
\footnotetext{\emph{2010 Mathematics Subject Classification.}
Primary: 46E15, secondary: 30D15; 30H20; 46B45; 46E05. }%
\footnotetext{\emph{Key words and phrases.} Weighted Banach spaces of entire functions; Taylor coefficients; solid hull; solid core. }%

\section{Introduction and first results.} \label{sec1}

The aim of this paper is to investigate the solid hull of weighted Banach spaces  $H_v^\infty(\C)$ of all
entire functions $f$ such that $\Vert f \Vert_v:= \sup_{z \in \C} v(z)|f(z)|$ is finite. In what follows, we identify an entire function $f(z)=\sum_{n=0}^\infty a_n z^n$ with the sequence of its Taylor coefficients $(a_n)_{n=0}^\infty$. For example in the case $v(z) = e^{-|z|}, z \in \C,$ we show in
Theorem \ref{solidhull1} that the solid hull
consists precisely of complex sequences $(b_m)_{m=0}^\infty$ such that
\begin{equation}
\sup_{n \in \mathbb{N}}
\sum_{m=n^2 + 1 }^{(n+1)^2} |b_m|^2
e^{-2n^2} n^{4m}  
< \infty . \nonumber
\end{equation}
We are also able to characterize in Theorem \ref{solidhull}, the solid hulls for a quite general class of  weights in terms of numerical sequences defined by  Lusky, \cite{L},
in his investigations of the isomorphic classes of the spaces  $H_v^\infty(\C)$.
This class of weights includes those satisfying condition (B) of \cite{L}, see
Remark \ref{remLB} and Corollary \ref{corex}.
The calculation of the numerical sequences for some important weights $v$ is one of the
results of our paper, see Proposition \ref{exampleexponential}.
In addition to techniques of \cite{L}, our approach  uses
the methods of Bennet, Stegenga and Timoney in their paper \cite{BST}, where
the solid hull and the solid core of the weighted spaces $H^\infty_v(\D)$ were
determined for doubling weights $v$ on the open unit disc $\D$. In Section \ref{sec4} we show that our results in Section \ref{solid} can be used to determine space of multipliers from $H_v^\infty(\C)$ into $\ell^p, 1 \leq p \leq \infty$.

The solid hull and multipliers on spaces of analytic functions on the disc has been investigated by many authors. In addition to  \cite{BST}, we mention here a non exhaustive sample:  \cite{AS}, \cite{Bl}, \cite{BRV}, \cite{D}, \cite{JP}, \cite{V} and the list of references in \cite{BP}. Moreover, the papers \cite{BG}, \cite{Li} and \cite{Tung} investigate the behavior of the Taylor coefficients of entire functions belonging to weighted spaces similar to those considered in this paper.
Spaces of type $H_v^\infty(\C)$ and $H^\infty_v(\D)$ appear
in the study of growth conditions of
analytic functions and have been investigated in various articles since
the work of Shields and Williams, see {\it e.g.} \cite{BBG},\cite{BBT}, \cite{BT}, \cite{L0},
\cite{L}, \cite{SW} and the references therein.

A {\it weight} $v$   is a continuous function  $v: [0, \infty[ \to ]0,  \infty [$, which is non-increasing on $[0,\infty[$ and it is rapidly decreasing, i.e.\ it satisfies
$\lim_{r \rightarrow \infty} r^m v(r)=0$ for each $m \in \N$. We extend $v$ to $\C$ by $v(z):= v(|z|)$. For
such a weight, the {\it weighted Banach space of entire functions}
is defined by
\begin{center}
$H^\infty_v(\C) := \{ f \in H(\C) \, : \,  \Vert f \Vert_v := \sup_{z \in \C} v(|z|) |f(z)| <
 \infty \}$,
\end{center}
and it is  endowed with the weighted sup norm $\Vert \cdot  \Vert_v .$  Spaces of this type are also called sometimes \textit{weighted Fock spaces of infinite order}. For an entire function $f \in H(\C)$, we denote $M(f,r):= \max\{|f(z)| \ | \ |z|=r\}$. Using the notation $O$ and $o$ of Landau, $f \in H_v^\infty(\C)$ if and only if $M(f,r)=O(1/v(r)), r \rightarrow \infty$. The symbol $\N$ stands for the natural numbers $n=1,2,3,...$.

As we already mentioned, an entire function $f(z)=\sum_{n=0}^\infty a_n z^n$ is identified with the sequence of its Taylor coefficients $(a_n)_{n=0}^\infty$, that will be also denoted sometimes by $(a_n)_n$.
As is well-known, it is often impossible to characterize standard Banach spaces of
entire functions in terms of the Taylor coefficients; this is for example true
for the function spaces $H_v^\infty (\C)$.  The next best thing is to find the
strongest growth condition that the coefficients have to satisfy. This motivates
the concept of a solid hull, and we now recall the related
definitions and facts from \cite{AS}.

Let $A$ and $B$ be vector spaces of complex sequences containing the space of all the sequences with finitely many non-zero coordinates. The space $A$ is \textit{solid} if $a=(a_n) \in A$ and $|b_n| \leq |a_n|$ for each $n$ implies $b=(b_n) \in A$. The \textit{solid hull of $A$} is

$$S(A):= \{ (c_n) \, : \, \exists (a_n) \in A \ \mbox{such that} \ |c_n| \leq |a_n| \ \forall n \in \N \}. $$

\noindent The \textit{solid core of $A$} is

$$s(A):=\{ (c_n) \, : \, (c_na_n) \in A \ \forall (a_n) \in \ell_\infty \}.$$

\noindent The \textit{set of multipliers form $A$ into $B$} is

$$(A,B):= \{ c=(c_n) \, : \, (c_na_n) \in B \ \forall (a_n) \in A \}. $$

\medskip

\noindent \textbf{Facts:} 1. $A$ is solid if and only if $\ell_\infty \subset (A,A)$.

\noindent 2. $A \subset (B,C)$ if and only if $B \subset (A,C)$.

\noindent 3. The solid core $s(A)$ of $A$ is the largest solid space contained in $A$. Moreover $s(A)=(\ell_\infty,A)$.

\noindent 4. The solid hull $S(A)$ of $A$ is the smallest solid space containing $A$.

\noindent 5. If $X$ is solid, $(A,X)=(S(A),X)$ and $(X,A)=(X,s(A))$.

\medskip

We conclude this section with our first results.

\begin{proposition}\label{core}
The solid core of $H^\infty_v(\C)$ is
$$
s(v,\C):=\{ (a_n)_n \, : \, \Vert(a_n)\Vert_{1,v,\C} := \sup_{r>0} v(r) \sum_{n=0}^\infty |a_n| r^n < \infty \}.
$$
\end{proposition}
\begin{proof}
Given $(a_n)_n \in s(v,\C)$, the function $f(z)= \sum_{n=0}^\infty a_n z^n$ is clearly entire.
Moreover $v(|z|)|f(z)| \leq \Vert(a_n)_n\Vert_{1,v,\C}$ for each $z \in \C$, and $f \in H^\infty_v(\C)$.

To see the other inclusion, let $A$ be a solid sequence space contained in $H^\infty_v(\C)$, i.e.\ for each $(a_n)_n \in A$, $(|a_n|)_n \in A$ and $g(z)= \sum_{n=0}^\infty |a_n| z^n \in H^\infty_v(\C)$. Clearly, $M(g,r)= \sup_{|z|=r} \big| \sum_{n=0}^\infty |a_n| z^n \big| = \sum_{n=0}^\infty |a_n| r^n$. Therefore
$$\sup_{r>0} v(r) \sum_{n=0}^\infty |a_n| r^n = \sup_{r>0} v(r)M(g,r) < \infty.$$
\end{proof}

\begin{remark} Write for $n \in \N$, $\Vert z^n \Vert_v:= \sup_{r>0} v(r)r^n$. Clearly
$\Vert z^0\Vert_v=v(0)$. The weighted $\ell_1$ space $\ell_1((\Vert z^n\Vert_v)$ of all those complex sequences $(a_n)_n$ such that $\sum_{n=0}^\infty |a_n| \Vert z^n\Vert_v < \infty$ is contained in the solid core $s(v,\C)$. However, in general the inclusion is strict. Indeed, if $v(r)=e^{-r}, r \geq 0$, then $\Vert z^n\Vert_v = (n/e)^n$ as a direct calculation shows. Take $a_n:=1/n!, \ n=0,1,2,...$. For each $r>0$, we have $v(r) \sum_{n=0}^\infty  a_n r^n =1$. However, the series $\sum_{n=0}^\infty (n^n/n! e^n)$ diverges by the Stirling's formula $n! \sim (2 \pi n)^{1/2} (n/e)^n$.
\end{remark}

Our next elementary result about the behavior of the Taylor coefficients of elements $f \in H_v^\infty(\C)$, that holds for arbitrary weights $v$,  clarifies the importance of the study of the solid hull of $H_v^\infty(\C)$ in Section \ref{solid}.

\begin{proposition}\label{hull}
The Taylor coefficients of an entire function $f(z)=\sum_{n=0}^\infty a_n z^n$ in $H^\infty_v(\C)$ satisfy

\begin{itemize}

\item[(i)] $\sup_{r>0} v(r) \left(\sum_{n=0}^\infty |a_n|^2 r^{2n} \right)^{1/2} \leq \Vert f\Vert_v$, and

\item[(ii)] $\sup_{n} |a_n| \Vert z^n\Vert_v \leq \Vert f\Vert_v$.

\end{itemize}
\end{proposition}
\begin{proof}
$(i)$ Since $H^\infty$ is contained in $H^2$, for each function $h(z)= \sum_{n=0}^\infty b_n z^n$ analytic in a neighbourhood of $\D$, we have $(\sum_{n=0}^\infty |b_n|^2)^{1/2} \leq M(h,1)$. Now, given
$f(z)=\sum_{n=0}^\infty a_n z^n \in H^\infty_v(\C)$ and $r>0$, set $g(\zeta):=f(r \zeta)$, that is clearly an analytic function in a neighbourhood of $\D$. Since $g(\zeta)=f(r \zeta) = \sum_{n=0}^\infty a_n r^n \zeta^n$, we conclude
$$
\left(\sum_{n=0}^\infty |a_n|^2 r^{2n} \right)^{1/2} \leq M(g,1) =M(f,r).
$$
This implies the inequality in the statement (i).

(ii) follows from (i).
\end{proof}

\begin{remark} In general it is not true that $f \in H^\infty_v(\C)$ implies that $(|a_n| \Vert z^n\Vert_v)_n \in \ell_2$. To see this, take again $v(r)=e^{-r}, r \geq 0$. Clearly $e^z = \sum_{n=0}^\infty (1/n!) z^n$ belongs to $H^\infty_v(\C)$. However, the series $\sum_{n=0}^\infty (|a_n| \Vert z^n\Vert_v)^2 = \sum_{n=0}^\infty (n^n/n! e^n)^2$ diverges because $(n^n/n! e^n)^2 \sim 1/n$ by Stirling's formula.
\end{remark}

\section{The solid hull of $H^\infty_v(\C)$.}\label{solid}

We fix for this section a weight $v: \C \to ]0,\infty[$ satisfying the general hypothesis
made in Section \ref{sec1}. Our next aim is to   characterize  the solid hull of $H^\infty_v(\C)$ for weights
satisfying the additional condition \eqref{35}, below.
Let us start by introducing some notation used in \cite{L}.

We denote by $[x]$ be the largest
integer less or equal $x$ for a given real number $x \in \R$.
Given $m>0$, we denote by $r_m$ the global maximum point of $r^m v(r)$. Then $r_m \rightarrow \infty$ as $m \rightarrow \infty$. For example, if $v(r)= \exp(-\alpha r^p)$, then $r_m=(m/\alpha p)^{1/p}, \ m>0$.
Given an entire function $f(z)= \sum_{k=0}^\infty a_k z^k$, and $0<m<n$ (not necessarily integers) we define the following operators of de la Vall\'ee-Poussin type:
$$
V_{n,m}f:= \sum_{0 \leq k \leq m } a_k z^k + \sum_{m < k \leq n} \frac{[n]-k}{[n]-[m]} a_k z^k,
$$
$$
V_{p,0}f:= \sum_{0 \leq k \leq p}  \frac{[p]-k}{[p]} a_k z^k \, ;
$$
here and later,  the summation is performed over integers belonging to
the given intervals, although the endpoints of the intervals need not be
integers. We also denote, for $0<m<n$,
\begin{equation}
A(m,n):= \left(\frac{r_{m}}{r_{n}}\right)^{m} \frac{v(r_{m})}{v(r_{n})} \ \
\ \mbox{and}  \ \ \
B(m,n):= \left(\frac{r_{n}}{r_{m}}\right)^{n} \frac{v(r_{n})}{v(r_{m})}.
\label{28}
\end{equation}

Several results of Lusky, \cite{L}, will be needed below.
We start with the following lemma.

\begin{lemma} \label{luskylemma1}
(\cite{L}, Corollary 3.2 (b))
Let $0<m<n$ and let $Q(z)=\sum_{m < k \leq n} a_k z^k$ be a polynomial. Then
$$ \Vert Q\Vert_v \leq 2 A(m,n) \sup_{|z|=r_n} |Q(z)| v(z),$$
and
$$ \Vert Q\Vert_v \leq 2 B(m,n) \sup_{|z|=r_m} |Q(z)| v(z).$$

\end{lemma}

Given a strictly increasing sequence $(m_n)_{n=1}^{\infty}$ with $\lim_{n \rightarrow \infty} m_n =
\infty$, we define $m_0=0$ and $V_n := V_{m_{n+1},m_n} - V_{m_{n},m_{n-1}},
n \in \N$. For each $n \in \N, n \geq 2,$
$V_nf$ is a polynomial with all terms of degree  at least $m_{n-1} + 1 $ and
at most $m_{n+1}$, for all $f \in H(\C)$. In fact, for each $f(z)=\sum_{n=0}^\infty f_n z^n\in H(\C)$, we have
$$
V_nf(z)= \sum_{m_{n-1} < m \leq m_{n+1}} \gamma_m f_m z^m
= \sum_{m= [m_{n-1}] + 1}^{[ m_{n+1}]} \gamma_m f_m z^m ,
$$
where the numbers $\gamma_m \in [0,1]$  are
$$
\gamma_m = \frac{[m_{n+1}]-m}{[m_{n+1}]-[m_n]}, \ \ m_n < m \leq m_{n+1}
$$
and
$$
\gamma_m = \frac{m-[m_{n-1}]}{[m_{n}]-[m_{n-1}]}, \ \ m_{n-1} < m \leq m_{n}.
$$
The sum in $V_1f$ is understood to go from $m=0$ to $m=m_1$.

\begin{lemma}\label{luskylemma2} (\cite{L}, Prop. 3.4 (b))
If the sequence $(m_n)_n$ with $\lim_{n \rightarrow \infty} m_n = \infty$ satisfies
$$
\min(A(m_n,m_{n+1}),B(m_n,m_{n+1})) \geq b
$$
for some $b>2$, then there is $D>0$ such that $\Vert V_n\Vert  \leq D$ for each $n \in \N$, where $\Vert .\Vert $ is the operator norm in $L(H^\infty_v(\C))$ with respect to $\Vert \cdot \Vert_v$.
\end{lemma}

\begin{lemma} \label{luskylemma3} (\cite{L}, Lemma 5.1)
Fix $b>1$. For each weight $v: [0, \infty[ \rightarrow ]0, \infty[$ there is a sequence of numbers $0<m_1 <m_2<...$ with $\lim_{n \rightarrow \infty} m_n = \infty$, such that
$$
A(m_n,m_{n+1}) \geq b \ \ \ {\rm and} \ \ \ B(m_n,m_{n+1}) \geq b
$$
and
$$
\sup\limits_{n \in \N} \min(A(m_n,m_{n+1}),B(m_n,m_{n+1})) < \infty.
$$
\end{lemma}

In fact, Lusky proves in \cite[Lemma 5.1]{L} that $(m_n)_n$ can be selected to satisfy
$\min(A(m_n,m_{n+1}),B(m_n,m_{n+1}))=b$ for each $n \in \N$.
Our next lemma is contained in  Proposition 5.2 of \cite{L}, and its proof in \cite{L}
only uses the fact that $\min(A(m_n,m_{n+1}),B(m_n,m_{n+1})) \geq b$, not that
$\min(A(m_n,m_{n+1}),$ $B(m_n,m_{n+1}))=b$ for each $n \in \N$.
(In Remark \ref{rem9} we will be  able to clarify, which of these quantities $A$ or $B$
is larger in the case of the weights $\exp(-ar^p)$.)
Accordingly, the following result holds true, and it contains an important expression for a
norm equivalent to that of $H_v^\infty(\C)$.

\begin{lemma} \label{luskylemma4} (\cite{L}, Proposition 5.2)
Let the numbers $(m_n)_{n=1}^\infty$ and $b$ be as in Lemma
\ref{luskylemma3}, and in addition assume that  $b > 2$. Then, for every
$f \in H_v^\infty(\C)$,
\begin{eqnarray}
& & c_1 \sup\limits_{n \in \N} \sup\limits_{r_{m_{n-1}} \leq |z| \leq r_{m_{n+1}}}
|f_n(z)| v(z)
\leq
\Vert f \Vert_v \leq
c_2 \sup\limits_{n \in \N} \sup\limits_{r_{m_{n-1}} \leq |z| \leq r_{m_{n+1}}}
|f_n(z)| v(z) , \nonumber
\end{eqnarray}
where $f_n := V_n f$.

\end{lemma}

The main result of this section reads as follows.

\begin{theorem}
\label{solidhull}
Let $v : \mathbb{C} \to ]0,\infty[$ be a   weight
(which is continuous,  radial, non-increasing on $[0,\infty[$, and rapidly decreasing).
In addition, assume that there exists a sequence $0<m_1 <m_2<...$ with
$\lim_{n \rightarrow \infty} m_n = \infty$  such that for some $b>2$ and some $K \geq b$ we have
\begin{equation}
b \leq \min(A(m_n,m_{n+1}),B(m_n,m_{n+1})) \leq \max(A(m_n,m_{n+1}),B(m_n,m_{n+1})) \leq K,
\label{35}
\end{equation}
for each $n \in \N$. Then, the solid hull of $H^\infty_v(\C)$ is
\begin{equation}
S(v,\C):= \Big\{ (b_m)_{m=0}^\infty \, : \,
\sup_{n \in \N} v(r_{m_{n}})
\Big( \sum_{m_n < m \leq m_{n+1}} |b_m|^2
 r_{m_n}^{2m} \Big)^{1/2} < \infty \  \Big\} .  \label{37}
\end{equation}
\end{theorem}

\begin{remark} Notice that if a weight $v$, as in Theorem \ref{solidhull}, satisfies \eqref{35}, then
the conclusion of Lemma \ref{luskylemma4} also holds.  In fact,
many important weights do satisfy 
\eqref{35}, see Corollary \ref{corex} and Remark \ref{remLB}. Moreover, examples
with explicit calculations of the sequence $(m_n)_n$ will be presented in
Section 3.

Of course, there is no need to extend the condition \eqref{37} to the finitely many coefficients
$b_m$, $0 \leq m \leq m_1$.

In the characterization \eqref{37} we could as well replace
$r_{m_n}$  by $r_{m_{n+1}}$. This follows from the proof below, or as well from
the the right hand side inequality in the assumption \eqref{35} together with the definition \eqref{28}.
Namely, the condition \eqref{35} implies
\begin{equation}
\frac{1}{K} \leq \Big( \frac{r_{m_{n+1}}}{r_{m_n}} \Big)^{m_n}
\frac{v (r_{m_{n+1}})}{v(r_{m_n})}
\leq \Big( \frac{r_{m_{n+1}}}{r_{m_n}} \Big)^{m_{n+1}}
\frac{v (r_{m_{n+1}})}{v(r_{m_n})} \leq K.
\end{equation}
Hence, for every $m$ with $m_n \leq m \leq m_{n+1}$ we have
\begin{equation}
\frac{1}{K} \leq \Big( \frac{r_{m_{n+1}}}{r_{m_n}} \Big)^{m}
\frac{v (r_{m_{n+1}})}{v(r_{m_n})}  \leq K,
\end{equation}
or, for all $m_n \leq m \leq m_{n+1}$,
\begin{equation}
\frac{1}{K} r_{m_n}^{m} v(r_{m_n}) \leq r_{m_{n+1}}^{m} v (r_{m_{n+1}})
  \leq K r_{m_n}^{m} v(r_{m_n}).
\end{equation}
This means that $ r_{m_n}^{m} v(r_{m_n})$ may be replaced by
$ r_{m_{n+1}}^{m} v (r_{m_{n+1}})$ in \eqref{37}.
\end{remark}

\begin{proof}
The proof will be obtained in two steps.

\noindent
\textbf{Step 1.} \textit{If $g(z) = \sum_{m=0}^\infty b_m z^m\in H_v^\infty(\C)$, and $\Vert g\Vert_v \leq 1$, then there is $C>0$ such that
$$
\Big( \sum_{m_n < m \leq  m_{n+1}} |b_m|^2
 r_{m_n}^{2m} \Big)^{1/2}
\leq  \frac{C}{v(r_{m_{n}})}
$$
for all $ n \in\N$.}

This step is the analogue in our setting of Theorem 1.8 in \cite{BST}.

We estimate first the sum from $m_n$ to $(m_n + m_{n+1})/2$. To do this, observe that if
$m_n \leq m \leq (m_n + m_{n+1})/2$, then $\gamma_m \geq 1/2$.
We have
\begin{eqnarray}
& & \Big( \sum_{m=[m_n]+1}^{[(m_n + m_{n+1})/2]} |b_m|^2  r_{m_n}^{2m}  \Big)^{1/2}
\leq
2 \Big( \sum_{m=[m_n]+1}^{[(m_n + m_{n+1})/2]}
\gamma_m^2 |b_m|^2  r_{m_n}^{2m}  \Big)^{1/2}
\nonumber \\
& \leq &
2 \Big( \sum_{m= [m_{n-1}] +1}^{[m_{n+1}]} \gamma_m^2 |b_m|^2  r_{m_n}^{2m}  \Big)^{1/2}
= 2 \Big( \int\limits_{\partial \D}
  \Big| \sum_{ m_{n-1} < m \leq  m_{n+1} } \gamma_m b_m  r_{m_n}^{m} z^m \Big|^2 dz
 \Big)^{1/2}
\nonumber \\
& \leq &
 4 \pi  \sup\limits_{|z| = 1}
\Big| \sum_{m_{n-1} < m \leq  m_{n+1} } \gamma_m b_m  r_{m_n}^{m} z^m \Big|
= 4 \pi  \sup\limits_{|z| = r_{m_n} }
\Big| \sum_{m_{n-1} < m \leq  m_{n+1} } \gamma_m b_m  r_{m_n}^{m}
\Big( \frac{z}{ r_{m_n}}  \Big)^m  \Big|
\nonumber \\
& = &
\frac{4 \pi}{v(r_{m_{n}})}  v(r_{m_{n}}) \sup\limits_{|z| = r_{m_n} }
\Big| \sum_{m_{n-1} < m \leq  m_{n+1}} \gamma_m b_m  z^m  \Big|
\nonumber \\
& = &   \frac{4 \pi}{v(r_{m_{n}})}
v(r_{m_{n}}) \sup\limits_{|z| = r_{m_n} }
\Big| V_n g (z)   \Big|
\leq   \frac{4 \pi \Vert V_n g \Vert_v }{v(r_{m_{n}})}
\leq \frac{4 \pi D}{v(r_{m_{n}})} ,
\end{eqnarray}
since the operators $V_n$ are uniformly bounded with respect to $\Vert \cdot  \Vert_v$ by $D>0$, by Lemma \ref{luskylemma2} and \eqref{35}.

Now we estimate the sum from $(m_n + m_{n+1})/2 $ to $m_{n+1}$.
Observe that
$$
V_{n+1}g(z)= \sum_{m_{n} < m \leq m_{n+2}} \tilde\gamma_m b_m z^m
= \sum_{m = [m_{n}] + 1 }^{[ m_{n+2}]} \tilde\gamma_m b_m z^m,
$$
where the numbers $\tilde\gamma_m \in [0,1]$  are
$$
\tilde\gamma_m = \frac{m-[m_{n}]}{[m_{n+1}]- [m_{n}]}, \ \ m_{n} < m \leq m_{n+1},
$$
which increase from $\tilde\gamma_{m_n+1}$ till $\tilde\gamma_{m_{n+1}}=1$. If $(m_n + m_{n+1})/2 < m \leq
m_{n+1}$, we have $\tilde\gamma_m \geq 1/2$. Thus, proceeding similarly as we did before, we get
\begin{eqnarray}
& & \Big( \sum_{m=[(m_n + m_{n+1})/2]+1}^{[m_{n+1}]} |b_m|^2  r_{m_n}^{2m}  \Big)^{1/2}
\leq
2 \Big( \sum_{m=[(m_n + m_{n+1})/2]+1}^{[m_{n+1}]}
 \tilde\gamma_m^2 |b_m|^2  r_{m_n}^{2m}  \Big)^{1/2}
\nonumber \\
& \leq &
2 \Big( \sum_{m=[m_{n}]+1}^{[m_{n+2}]} \tilde\gamma_m^2 |b_m|^2  r_{m_n}^{2m}  \Big)^{1/2}
\nonumber \\
& \leq&
\frac{4 \pi}{v(r_{m_{n}})}
v(r_{m_{n}}) \sup\limits_{|z| = r_{m_n} }
\Big| V_{n+1} g (z)   \Big|
\leq   \frac{4 \pi \Vert V_{n+1} g \Vert_v }{v(r_{m_{n}})}
\leq \frac{4 \pi D}{v(r_{m_{n}})}.
\end{eqnarray}
This completes the proof of Step 1. \\

Observe that the estimates proved in Step 1 remain valid if we replace $r_{m_n}$
by any  $r_m$ with $m_n \leq m \leq m_{n+1}$. \\

\noindent
\textbf{Step 2.} \textit{For each $(b_m)_{m=0}^\infty \in S(v,\C)$ there is  $f(z)= \sum_{m=0}^\infty a_m z^m  \in H_v^\infty(\C)$ such that $|b_m| \leq |a_m|$ for each $m=0,1,2,...$.}

Fix $C_0 > \sup_{n} v(r_{m_{n}})
\Big( \sum_{m_n < m \leq m_{n+1}} |b_m|^2
 r_{m_n}^{2m} \Big)^{1/2}$.

For all $n  \in\N$ and
$m_n <  m \leq m_{n+1}$,  we apply \cite{BST}, Corollary 2.8  (which is a consequence of a deep result of Kisliakov \cite{Ki}), to the sequence
$$
\big( b_{m  }  r_{m_n}^{m} \big)_{m_n < m \leq m_{n+1}  },
$$
to choose a polynomial
$$
P_n (z) = \sum_{m_n < m \leq m_{n+1} } b_m' z^m
$$
such that
$$
|b'_m| \geq |b_m|  r_{m_n}^{m} \ \forall m \  \ \ \  \mbox{and}
$$
\begin{equation}
\sup\limits_{|z| < 1 } | P_n (z)| = \sup\limits_{|z| = 1 } | P_n (z)|\leq B
\Big( \sum_{m_n < m \leq m_{n+1} } |b_m |^2  r_{m_n}^{2m} \Big)^{1/2}  .
\label{45}
\end{equation}
Here  $B>0$ is an absolute constant. Define
$$
Q_n (z) := \sum_{m_n < m \leq m_{n+1} }  r_{m_n}^{- m}  b_m' z^m
$$
and
$$
g(z) =  \sum_{n=1}^\infty  Q_n(z) =  \sum_{n=1}^\infty
\sum_{m_n < m \leq m_{n+1} } r_{m_n}^{-m}  b_m' z^m  .
$$
We still have to show that $g$ is a well defined entire function and that $g \in H_v^\infty(\C)$. However, observe that if $a_m$ denotes the $m$-th Taylor coefficient of $g$, then $|a_m| \geq |b_m|$ for all $m$.

\medskip

By Lemma \ref{luskylemma1}, \eqref{35} and \eqref{45}, we have
\begin{eqnarray}
& &
\Vert Q_n \Vert_v 
\leq 2K \sup\limits_{|z| = r_{m_n} } |  Q_n  (z) | v(z)
=  2K  \sup\limits_{|z| = r_{m_n} } \Big| \sum_{m_n < m \leq m_{n+1}}
r_{m_n}^{- m} b_m' z^m  \Big|
v(z)
\nonumber \\
&=&
2K \sup\limits_{|z| = r_{m_n} }  \Big| \sum_{m_n < m \leq m_{n+1} } b_m'
( z / r_{m_n})^m  \Big|  v(z)
=  2K v ( r_{m_n} ) \sup\limits_{|z| = 1}
\Big| \sum_{m_n < m \leq m_{n+1} } b_m'   z^m  \Big|
\nonumber \\
&=&
2K v ( r_{m_n} ) \sup\limits_{|z| < 1}
|P_n (z)|
\nonumber \\
&\leq &
2KB v( r_{m_n}) \Big( \sum_{m_n < m \leq m_{n+1}} r_{m_n}^{2m} |b_m|^2 \Big)^{1/2}
 \leq 2KBC_0, \nonumber
\end{eqnarray}
for each $n \in \N$.
Moreover, for all  $n \geq 2$, we have
\begin{eqnarray}
& &
\sup\limits_{r_{m_{n-1}} \leq |z| \leq r_{m_{n+1}}} |V_n  (Q_n + Q_{n-1} ) (z)|v(z)
\leq
\sup\limits_{z \in \C} |V_n  (Q_n + Q_{n-1} ) (z)|v(z)
\nonumber \\
&\leq &
D \sup\limits_{z \in \C}  |( Q_n + Q_{n-1})  (z)| v(z)
\leq D \Vert Q_n \Vert_v +  D \Vert Q_{n-1} \Vert_v \leq 4KBC_0D.
\label{47a}
\end{eqnarray}
Consequently for each $N >n \geq 2$, we have $V_n(\sum_{j=2}^N Q_j)(z)= V_n (Q_n + Q_{n-1})(z), z \in \C$.
By  Lemma  \ref{luskylemma4}     and \eqref{47a}
we get for every  $N \in \N$,
\begin{eqnarray}
& &
\Vert \sum_{j=2}^N Q_j \Vert_v = \sup\limits_{z \in \C} v(z)
\Big|\sum_{j=2}^N Q_j(z) \Big|
\nonumber \\
&\leq &
c_2 \sup\limits_{n \in \N} \sup\limits_{r_{m_{n-1}} \leq  \atop
|z| \leq r_{m_{n+1}}}
\Big|V_n \Big( \sum_{j=2}^N Q_j(z) \Big) \Big|v(z)
\nonumber \\
&= &
c_2 \sup\limits_{n \in \N}
 \sup\limits_{r_{m_{n-1}} \leq \atop |z| \leq r_{m_{n+1}}}
|V_n  (Q_n + Q_{n-1} ) (z)|v(z)
\leq
4KBC_0 c_2 D.
\end{eqnarray}
This implies that the sequence of polynomials $\left( \sum_{j=2}^N Q_j\right)_N$ is
contained in a multiple of the unit ball of $H_v^\infty(\C)$, which is compact for the
compact open topology. Accordingly, there is a subsequence $\left( \sum_{j=2}^{N(s)}
Q_j\right)_s$ converging to $h \in H_v^\infty(\C)$ for this topology. Since the operator of
$k$-th differentiation is continuous for the compact open topology, it follows that the
Taylor coefficients of $h$ and $g-Q_1$ coincide. This implies that $g$ is an entire function
and that $g \in H_v^\infty(\C)$. The proof of Step 2 is now complete by taking the function $f(z):=g(z)+\sum_{0 \leq m \leq m_1} b_m z^m$.
\end{proof}

\begin{remark}
\label{remLB}
(1) Lusky introduces the following \textit{condition (B)} on the weight $v$ in \cite{L}:
$$
\forall b_1>0 \ \exists b_2 >1 \ \exists c>0 \ \forall m,n :
$$
$$
\left(\frac{r_m}{r_n}\right)^m \frac{v(r_m)}{v(r_n)} \leq b_1 \ \ {\rm and} \ \ |m-n| \geq c \Rightarrow \left(\frac{r_n}{r_m}\right)^n \frac{v(r_n)}{v(r_m)} \leq b_2.
$$
By \cite{L}, Theorem 1.1, if $v$ has condition (B), then $H_v^\infty(\C)$ is isomorphic to $\ell_\infty$, and if $v$ does not satisfy condition (B), then $H_v^\infty(\C)$ is isomorphic to $H^\infty$.

We show that if the weight $v$ satisfies condition (B), then for each $b>2$ one can find a sequence $(m_n)_n$ satisfying the assumption (\ref{35}) in Theorem \ref{solidhull}. Indeed, let $v$ be a weight satisfying condition (B). Given $b>2$ we apply Lemma \ref{luskylemma3} to find a sequence $0<m_1 <m_2<...$ with $\lim_{n \rightarrow \infty} m_n = \infty$, such that
$A(m_n,m_{n+1}) \geq b, B(m_n,m_{n+1}) \geq b$ and $M:= \sup_n \min(A(m_n,m_{n+1}),B(m_n,m_{n+1})) < \infty$. An inspection of the proof of Lemma 5.1 in Lusky shows that we can take in our Lemma \ref{luskylemma3} the sequence $(m_n)_n$ such that $\lim_{n \rightarrow \infty} (m_{n+1}-m_n) = \infty$. Set $b_1 := M$, and select $b_2>1$ and $c>0$ according to condition (B). There is $n(0) \in \N$ such that $m_{n+1}-m_n \geq c$ if $n \geq n(0)$. Condition (B) now implies that $\max(A(m_n,m_{n+1}),B(m_n,m_{n+1})) \leq \max(M,b_2)$ for each $n \geq n(0)$. The proof is complete if we take $(m_n)_{n \geq n(0)}$. \\

(2) Lusky constructs in \cite{L}, Example 2.6, a weight $v$ on $\C$ not satisfying condition (B) such that for a certain sequence $(m_n)$ with $m_{n+1}-m_n = n+1$, $A(m_n,m_{n+1})=(n+1)^{n+1}$ and $B(m_n,m_{n+1})=1$ for each $n \in \N$.
\end{remark}

This remark and Examples 2.1--2.2 in \cite{L} imply the following result.

\begin{corollary}
\label{corex}
Condition \eqref{35} is satisfied by the following weights:

\noindent  $v(r)= \exp(-r^p)$ with  $p > 0$,
$v(r)= \exp( -\exp r)$,  and
$v(r)= \exp\big(- (\log^+ r  )^p\big) $, where $p \geq 2$ and $\log^+ r
= \max(\log r,0)  $.
\end{corollary}

\section{Examples.}

In this section we calculate the sequences $(m_n)_{n=1}^\infty$ for the weights
$v(r) = \exp ( -ar^p) $ and thus obtain satisfactory representations of the
corresponding solid hulls.

\begin{theorem}
\label{solidhull1}
Let $v $ be the weight $v(r) = \exp ( -ar^p) $ on $\mathbb{C}$, where $a> 0$ and $p>0$ are
constants. Then,  the solid hull of $H^\infty_v(\C)$ is
\begin{equation}
\Big\{ (b_m)_{m=0}^\infty \, : \,
\sup_{n \in \mathbb{N}}
 \sum_{pn^2 + 1 < m \leq p(n+1)^2} |b_m|^2
e^{ -2 n^2} n^{4m/p} (ap)^{-m/p}  < \infty \  \Big\} .  \label{55}
\end{equation}
\end{theorem}

In particular, the solid hull for $v(r)= \exp (-r) $ is
$$
\Big\{ (b_m)_{m=0}^\infty \, : \,
\sup_{n \in \mathbb{N}}
\sum_{m=n^2 + 1 }^{(n+1)^2} |b_m|^2
e^{-2n^2} n^{4m}  
< \infty \  \Big\} .
$$

Theorem \ref{solidhull1} is an immediate consequence of Theorem \ref{solidhull}
and the following proposition, where we choose $b = e$. The proposition gives
the Lusky numbers  $(m_n)_n$ for a class of important weights.

\begin{proposition}\label{exampleexponential}
Let $v(r)= \exp(-a r^p)$, $a>0$, $p>0$ and let $b>2$. The sequence $m_n:= p (\log b) n^2$,
$ n \in \N$, satisfies, for each $n \in \N$, $n\geq 4$,
$$
b \leq A(m_n,m_{n+1})= \left(\frac{r_{m_n}}{r_{m_{n+1}}}\right)^{m_n} \frac{v(r_{m_n})}{v(r_{m_{n+1}})} \leq b^{9/2}
$$
and
$$
b \leq B(m_n,m_{n+1})= \left(\frac{r_{m_{n+1}}}{r_{m_n}}\right)^{m_{n+1}} \frac{v(r_{m_{n+1}})}{v(r_{m_n})} \leq b^{4}.
$$
\end{proposition}

Proposition \ref{exampleexponential} implies that the numbers $m_n$
can chosen to be $n^2$ for the weight
$v(r) = \exp(-r)$ and $2 n^2$ for the weight $v(r) = \exp(-r^2)$.
We give the proof of this proposition in several steps.

\begin{lemma}\label{log1}
If $0<m<M$ satisfies $M<2m$, then
$$
\exp\left(\frac{1}{2}\frac{(M-m)^2}{M}\right) \leq \left(\frac{M}{m}\right)^M e^{m-M} \leq
\exp\left(\frac{(M-m)^2}{M}\right).
$$

\end{lemma}
\begin{proof} If $0<x<1/2$, then
\begin{equation}
 x^2 / 2 \leq - \log(1-x) -x \leq x^2 \label{65}
\end{equation}
This is so since
$$
-\log(1-x)-x= x^2/2 + x^3/3 +... \geq x^2/2
$$
and on the other hand,
$$
x^2+\log(1-x)+x= x^2/2 - x^3/3 -... \geq  0.
$$
We now set $x:=(M-m)/M$. Clearly $0<x<1/2$, $1-x=m/M$ and $m-M=-Mx$. Hence
$$
M(-\log(1-x)-x) = -M \log(m/M) + m - M
$$
and \eqref{65} implies $Mx^2 / 2 \leq - M\log(1-x) - Mx \leq Mx^2$, or,
$$
\frac{1}{2}\frac{(M-m)^2}{M} \leq  M \log(M/m) + m - M
\leq \frac{(M-m)^2}{M} . 
$$
\end{proof}

\begin{lemma}\label{log2}
If $0<m<M$ satisfies $M < 7m / 4$, then
$$
\exp\left(\frac{1}{4}\frac{(M-m)^2}{m}\right) \leq \left(\frac{m}{M}\right)^m e^{M-m} \leq
\exp\left(\frac{1}{2} \frac{(M-m)^2}{m}\right).
$$
\end{lemma}
\begin{proof}
We have for $0<x<3/4$
$$
\log(1+x)-x=-x^2/2 + x^3/3 - x^4/4 + ... \geq -x^2/2.
$$
and
$$
-x^2/4 -\log(1+x)+x = x^2/4 - x^3/3 + x^4/4 + ... \geq x^2(1/4- x/3) >0.
$$
Hence, for these $x$,
\begin{equation}
-x^2 / 2  \leq \log(1+x) -x \leq - x^2 / 4 . \label{70}
\end{equation}
Fix $0<m<M< 7m / 4$. Set $x:=M/m-1$ so that $0<x<3/4$. We have
$$
-m(\log(1+x)-x)= m \log(m/M) + M - m,
$$
hence, \eqref{70} implies
$$
\frac{m}{4} x^2 = \frac{1}{4}\frac{(M-m)^2}{m} \leq m \log(m/M) + M - m \leq \frac{1}{2} \frac{(M-m)^2}{m} = \frac{m}{2} x^2.
$$
\end{proof}

\begin{lemma}\label{estimates}
Assume that $m_n= \alpha n^2, n \geq 4,$ for some $\alpha >0$. Then
$$
\exp \alpha \leq \left(\frac{m_{n}}{m_{n+1}}\right)^{m_{n}} \exp(m_{n+1}-m_n) \leq \exp( 9\alpha/4),
$$
and
$$
\exp \alpha \leq
 \left(\frac{m_{n+1}}{m_n}\right)^{m_{n+1}} \exp(m_n - m_{n+1})
 \leq \exp(4 \alpha).
$$
\end{lemma}
\begin{proof}
Set, for $n \geq 4$, $m:=\alpha n^2$ and $M:=\alpha (n+1)^2$. Then $M-m=\alpha(2n+1)$ and $M < 7m/4 < 2m$. It is now easy to see that
$$
4 \alpha \leq \frac{(M-m)^2}{m} \leq 9 \alpha \ \ \mbox{and} \ \
2 \alpha \leq \frac{(M-m)^2}{M} \leq 4 \alpha,
$$
The conclusion follows from Lemmas \ref{log1} and \ref{log2}.

\end{proof}

\begin{proof} \textbf{of Proposition \ref{exampleexponential}.}
In this case the maximum point $r_m$ of $r^m v(r)$ is $r_m=(m/ap)^{1/p}$,  and $v(r_m)=\exp(-m/p)$ for each $m \in \N$. Therefore
$$
A(m_n,m_{n+1})^p = \left(\frac{m_{n}}{m_{n+1}}\right)^{m_{n}} \exp(m_{n+1}-m_n),
$$
and
$$
B(m_n,m_{n+1})^p = \left(\frac{m_{n+1}}{m_n}\right)^{m_{n+1}} \exp(m_n - m_{n+1}).
$$
We apply Lemma \ref{estimates} for $\alpha:=p (\log b)$ to conclude
$$
b^p=\exp(\alpha) \leq A(m_n,m_{n+1})^p \leq \exp(9\alpha / 4) = b^{9p/4} ,
$$
and
$$
b^p=\exp(\alpha) \leq B(m_n,m_{n+1})^p \leq \exp(4 \alpha) = b^{4p}.
$$
This implies the inequalities in the statement.
\end{proof}

\begin{remark} \label{rem9}
In view of Lemmas \ref{luskylemma1}--\ref{luskylemma3} it is of interest to compare
the expressions $A(m,n) $ and $B(m,n)$. Let us show here that
$$A(m_n,m_{n+1}) \leq B(m_n,m_{n+1}) \ \ (*)$$ for $v(r) = \exp(-ar^p)$ and
$m_n = \alpha n^2, \alpha >0$. It is enough to do the calculation for $m_n = n^2$. In this case we have
$$
A(m_n,m_{n+1})= \Big( \frac{n}{n+1}\Big)^{2n^2}e^{2n+1} \ \ , \ \
B(m_n,m_{n+1})=  \Big( \frac{n+1}{n}\Big)^{2(n+1)^2}e^{-2n-1},
$$
so that $(*)$ is equivalent to
$$
\frac{\exp(n+(1/2))}{\left(1 + \frac{1}{n}\right)^{n^2}} \leq \frac{\left(1 + \frac{1}{n}\right)^{(n+1)^2}}{\exp(n+(1/2))},
$$
or 
\begin{equation}
e \leq \Big(1 + \frac{1}{n}\Big)^{\frac{2n^2+2n+1}{2n+1}} =: \gamma_n.
\label{105}
\end{equation}
But $\gamma_n > \left(1 + \frac{1}{n}\right)^{n + \frac{1}{2}}=: \eta_n.$ Thus,
\eqref{105} holds true, since the sequence $(\eta_n)_n$
tends to $e$ as $n \rightarrow \infty$, and it is decreasing.
To see this last fact, we write 
\begin{equation}
\Big( \frac{\eta_n}{\eta_{n+1}} \Big)^2
= \frac{(n+1)^{4n+4}}{n^{2n+1} (n+2)^{2n+3}}
= \bigg(
\frac{n+1}{n^{\frac{2n+1}{4n+4}} (n+2)^{\frac{2n+3}{4n+4}}}
\label{107}
\bigg)^{4n+n}
\end{equation}
The logarithm of the expression in the last large parenthesis is of the form
$$
\log x - \big(r \log (x-1) + (1-r) \log (x+1) \big),
$$
where $x > 1$ and $0< r < 1$. This expression is positive, since log
is a concave function on $[0,\infty[$. Hence, \eqref{107} is larger than 1.
\end{remark}

\section{The space of multipliers $( H^\infty_v(\C), \ell^p ) $.}\label{sec4}

In this final section we show that the above results can be applied to
determine some multiplier spaces. Recall that if $A$ and $B$ are vector spaces of complex sequences containing the space of all the sequences with finitely many non-zero coordinates, then the set of multipliers from $A$ into $B$ is

$$(A,B):= \{ c=(c_n) \, : \, (c_na_n) \in B \ \forall (a_n) \in A \}. $$

Given a strictly increasing, unbounded sequence $J= (m_n)_{n=0}^\infty
\subset \N$  and  $1 \leq p,q \leq \infty$ we denote as in \cite{BlZ}, Definition 2,
\begin{equation}
\ell^J (p,q) := \Big\{ (a_m)_{m=0}^\infty \, : \, \Big( \sum_{m=m_n+1}^{m_{n+1}}
|a_m|^p \Big)^{1/p} \in \ell_q \Big\},
\end{equation}
with the obvious changes when $p$ or $q$ is $\infty$. The space $\ell^J (p,q)$
is a Banach space when endowed with  the canonically defined norm. Observe
that $\ell^J(p,p) = \ell_p$.

\begin{lemma}
\label{lem5.1}
For $1 \leq p \leq \infty$ we have
\begin{equation}
\big( \ell^J(2,\infty) , \ell^p \big) =
\ell^J(r,s)
\end{equation}
where (a) $r=2p/(2-p)$, $s=p$, if $1 \leq p < 2$, (b) $r=\infty$,
$s=p$, if $2 \leq p < \infty$, and (c) $r= s=\infty$, if $p = \infty$.
\end{lemma}

\begin{proof}
This is a direct consequence of \cite{BlZ}, Theorem 23. As it is mentioned in that
paper, which treats more complicated cases, the proof in the case of our Lemma is similar to that of \cite{K}, Theorem 1.
\end{proof}

\begin{theorem}
\label{th5.2}
Let $v$ be a radial weight and let  $J= (m_n)_{n=0}^\infty $ be a strictly increasing, unbounded sequence of positive integers such that for some $b > 2$ and $K \geq b$ we have
\begin{equation}
b \leq \min(A(m_n,m_{n+1}),B(m_n,m_{n+1})) \leq \max(A(m_n,m_{n+1}),
B(m_n,m_{n+1})) \leq K.
\label{135}
\end{equation}
Let $ 1 \leq p \leq  \infty$. Then $(\lambda_m)_{m=1}^\infty$ is a multiplier
from $H_v^\infty(\C)$ into $\ell_p$ if and only if
\begin{equation}
\Big( \big( (v(r_{m_n}) r_{m_n}^m )^{-1} |\lambda_m| \big)_{m=m_n+1}^{m_{n+1}}
\Big)_{n=0}^\infty \in \ell^J(r,s) ,
\end{equation}
where

\noindent (a) $r=2p/(2-p)$, $s=p$, if $1 \leq p < 2$,

\noindent (b) $r=\infty$, $s=p$, if $2 \leq p < \infty$, and

\noindent (c) $r= s=\infty$, if $p = \infty$.

\end{theorem}

\begin{proof}
Since $\ell^p$ is a solid space, we can apply Fact 5 in the introduction to
conclude
\begin{equation}
\big( H_v^\infty( \C) , \ell_p \big)=
\big( S( H_v^\infty( \C)) , \ell_p \big) .
\end{equation}
By Theorem \ref{solidhull} it is easy to see that
$(\lambda_m)_{m=0}^\infty \in \big(  S( H_v^\infty( \C)) , \ell_p \big)$, if and
only if
\begin{equation}
\Big( \big( (v(r_{m_n}) r_{m_n}^m )^{-1} |\lambda_m| \big)_{m=m_n+1}^{m_{n+1}}
\Big)_{n=0}^\infty \in \big( \ell^J(2, \infty) , \ell_p\big) .
\end{equation}
The conclusion now follows from Lemma \ref{lem5.1}
\end{proof}

The next corollary is a consequence of Theorem \ref{th5.2} and
Proposition \ref{exampleexponential}.

\begin{corollary}
Let $v(r) = e^{-r}$, $r \in (0,\infty)$ and $1 \leq p \leq \infty$.
Then, the space of multipliers $\big( H_v^\infty(\C), \ell_p \big)$ is the set of sequences
$(\lambda_m)_{m=0}^\infty$ such that
\begin{equation}
\Big( \sum_{n=1}^\infty  \Big( \sum_{m=n^2 +1}^{(n+1)^2}
\big( |\lambda_m| e^{-n^2} n^{-2m}  \big)^{\frac{2p}{2-p}}
\Big)^{\frac{2-p}{2}} \Big)^{\frac{1}{p}} < \infty  ,
\end{equation}
if $1 \leq p < 2$,
\begin{equation}
\Big( \sum_{n=1}^\infty  \Big( \max\limits_{n^2 < m \leq (n+1)^2}
 |\lambda_m| e^{n^2} n^{-2m}  \Big)^{p} \Big)^{\frac{1}{p}} < \infty  ,
\end{equation}
if $2 \leq p < \infty$, and
\begin{equation}
\sup\limits_{n \in \N } \Big( \max\limits_{n^2 < m \leq (n+1)^2}
 |\lambda_m| e^{n^2} n^{-2m} \Big) < \infty  ,
\end{equation}
if $p =  \infty$,

\end{corollary}

\vspace{.5cm}

\noindent \textbf{Acknowledgements.} The authors are indebted to the referee for the careful reading of the manuscript and the helpful suggestions.

The research of Bonet was partially supported by MTM2013-43540-P, GVA Prometeo II/2013/013 and GVA ACOMP/2015/186. The research of Taskinen was partially supported by the Magnus Ehrnrooth and the V\"ais\"al\"a
Foundations.

\noindent \textbf{Authors' addresses:}%
\vspace{\baselineskip}%

Jos\'e Bonet: Instituto Universitario de Matem\'{a}tica Pura y Aplicada IUMPA,
Universitat Polit\`{e}cnica de Val\`{e}ncia,  E-46071 Valencia, Spain

email: jbonet@mat.upv.es \\

Jari Taskinen: Department of Mathematics and Statistics, P.O. Box 68,
University of Helsinki, 00014 Helsinki, Finland.

email: jari.taskinen@helsinki.fi

\end{document}